Coronado Industries had January 1 inventory of $301000 when it adopted dollar-value LIFO. During the year, purchases were $1780000 and sales were $3050000. December 31 inventory at year-end prices was $427350, and the price index was 111.

What is Coronado Industries’s ending inventory?
Answer #1

Ending Inventory = $427350/1.11 = $388,500

Ending Inventory = $388,500 - 301000 = $87,500

Ending Inventory = 301000 + 87,500*1.11 = $397,250